OpenStudy (winsteria):

Solve for n. 9/8n = -54

OpenStudy (winsteria):

@abswag13

OpenStudy (anonymous):

n = -48

OpenStudy (winsteria):

Okay thank you

OpenStudy (skullpatrol):

$$\Huge \frac{9}{8} n = -54$$

OpenStudy (skullpatrol):

multiply both sides by 8/9

OpenStudy (britanyashton):

multiply each side of the equation by 8 9/8n=-54 *8 *8 9n=-432 /9 /9 n=-48